OpenStudy (anonymous):

Write the equation of a circle with a center at (-2, -5) and a radius of 6.5. THE answer is (x + 2)2 + (y + 5)2 = 42.25?

OpenStudy (anonymous):

(x + 2)^2 + (y + 5)^2 = 42.25?

OpenStudy (apoorvk):

a circle with a centre (h,K) and radius "r" has an equation: (x-h)^2 + (y-k)^2 = r^2 you are right!

OpenStudy (anonymous):

Write the equation of a circle with a center at (-3, -8) and a radius of 1/3 the answer is (x + 3)^2 + (y + 8)^2 = 1/3 ?

OpenStudy (apoorvk):

no the constant on RHS will be 1/9... coz you square 1/3.. rest is right

OpenStudy (anonymous):

I have one last question that can you check ?

OpenStudy (anonymous):

What is the point of symmetry for the following circle? http://i729.photobucket.com/albums/ww292/akatushi101/mamita.gif

OpenStudy (apoorvk):

i think point of symmetry means the centre which should be (-3,-1).

OpenStudy (anonymous):

Thanks for all the help :)